OpenStudy (anonymous):

Given the function g(x) = 8x − 2, find g(−2).

OpenStudy (anonymous):

Plug x = -2, ...

OpenStudy (anonymous):

I think I got the answer, is it -18?

OpenStudy (anonymous):

Correct :) g(-2) = 8(-2) - 2 = -16 - 2 = -18 @vetdoc I'll expect next time you explain your result like I showed you this time.

OpenStudy (anonymous):

ooh yay!! I did it like that!! :D

OpenStudy (anonymous):

Remember to show your work, so the helpers understand how your brain processes :)

OpenStudy (anonymous):

okay, and I have one really quick and easy question.

OpenStudy (anonymous):

Can I ask here really quick? :)

OpenStudy (anonymous):

Yes, if it's "really really quick" ;)

OpenStudy (anonymous):

The formula for degrees Celsius is: C = five over nine (F − 32), where F stands for degrees Fahrenheit. Which of the following equations shows the formula correctly solved for F? F = five over nine C + 32 F = nine over five C + 32 F = five over nine C − 32 F = nine over five C − 32

OpenStudy (anonymous):

c = 5/9 ( F - 32 ) F = 9/5c + 32

OpenStudy (anonymous):

just make F subject of formula i guess...

OpenStudy (anonymous):

Can you expand from the parentheses?

OpenStudy (anonymous):

@vetdoc Do you understand what antoni7 did?